A box containing 3 oranges, 2 apples, and one banana weighs 15 units. Another box containing 5 oranges, 7 apples, and 2 bananas weighs 44 units. A third box containing 1 orange, 3 apples, and 5 bananas weighs 26 units. How much does each fruit weigh?
Please help.

Answer by flame8855(424)

let x be the weight of oranges
let y be the weight of apples
let z be the weight of banana
3x+2y+z= 15
5x+7y+2z = 44
x+3y+5z=26
solve the three equations
x= 1 unit (orange)
y = 5 unit (apple)
z= 2 unit (banana)